A 100% pure sample of a divalent metal carbonate weighing 2g on complete thermal decomposition released 448 cc of co2 at STP .the equivalent weight of the metal?

Grade:11

1 Answers

Sonali Purkait
5 years ago
Divalent metal carbonate means MCO3MCO3 ------->. MO + CO2 1 mol. 22400 cc 2 g. 448 cc 448 cc of CO2 is evolved from 2g of MCO3 22400 cc of CO2 will be evolved from ( 2× 22400)÷448 = 100g of MCO3i.e. 100g is the molecular wt. of the carbonate Then mol. wt. of metal = 100 - (12+3×16){wt. of CO3 2- }=40gEquivalent wt. = Molecular wt. / Valency = 40/2=20